OpenStudy (anonymous):

An architect needs to determine the slope between two points on a ski lift. The two points have been identified as (25, 30) and (125, 50), where x is the horizontal distance and y is the vertical distance from the bottom of the lift. Assuming the lift runs in a straight line, what is the slope of the line between the two points? (Write your answer in simplest form, using / for a fraction bar if needed.) My answer is 1.5 Am I right? :)

OpenStudy (abb0t):

\[slope = m = \frac{ y_2-y_1 }{ x_2-x_1 }\] then, choose any of the two points and solve in the form y = mx+b where you will be using: \[(y-y_1)=m(x-x_1) OR(y-y_2) = m(x-x_2)\]